The problem is:
A man lives 12 miles from work and 7 miles from the grocery store. If he made x round trips to the store, write an expression to represent how many miles he drove? Show work for credit.

Answer by JulietG(1812)

His distance to work is irrelevant here. It's thrown in to confuse you.
Let x be the number of round trips to the store. Let y be the total distance traveled.
x * (7+7) [because it's round trip]=y
14x = y
.
So, if the man made 2 round trips, x would be 2.
14*2 = 28
If he made 22, x would be 22, and y would change accordingly.
14*22 = 308
etc